OpenStudy (anonymous):

Second-Generation Offspring In another experiment, Mendel crossbred non pure tall pea plants. As a result, the second-generation offspring were 787 tall plants and 277 short plants. Determine the empirical probability of a second-generation plant being.

OpenStudy (anonymous):

find P(tall plants) and P(short plants)

OpenStudy (anonymous):

i founds it!

OpenStudy (anonymous):

You serious?

OpenStudy (anonymous):

well your question at least....lol

OpenStudy (anonymous):

but i might have an answer for you

OpenStudy (anonymous):

Lol god finally

OpenStudy (anonymous):

You know why you get hit on because you got your pic up as an icon.

OpenStudy (anonymous):

okay so knowing nothing about biology or whatever youre taking, i looked up how to determine the empirical probability....heres what i found (for your exact question at that!) The cross is: Tt x Tt possible offspring TT Tt tT tt = 3/4 tall 1/4 short.

OpenStudy (anonymous):

well i am signed on through facebook. lol

OpenStudy (anonymous):

My professor didn't accept that answer

OpenStudy (anonymous):

And you googled that lmao

OpenStudy (anonymous):

hey i tried. lol thats how i try to figure out my problems. usually it helps. haha im trying for ya!

OpenStudy (anonymous):

True now do you want help?

OpenStudy (anonymous):

Maybe I could try to answer it

OpenStudy (anonymous):

please for the love of god. or ill jump out my window. only problem is im on the first floor. lmao

OpenStudy (anonymous):

Omg you biting my lines

OpenStudy (anonymous):

hey at least i was paying attention unlike everyone else in this place. haha

OpenStudy (anonymous):

True you tried gosh lol

OpenStudy (anonymous):

hold on im still looking for ya....

OpenStudy (anonymous):

Giv me your question in the mean while though

OpenStudy (anonymous):

im having trouble getting the answer to the integral of (1)/(49-169x^2). my answer does not match the real answer, and im not sure how they got their answer. please help! i need step by step instructions!!!

OpenStudy (anonymous):

What is full question, I mean can you post your whole question on here

OpenStudy (anonymous):

that was my question. lol

OpenStudy (anonymous):

Lol I can give you the ans but not step by step

OpenStudy (anonymous):

my answer is this....1/98 * (ln(7+13x)+ln(7-13x)) +c the real answer is ....1/182 * (ln(7+13x)-ln(7-13x)) + c i just dont understand why its 1/182 and where they got the minus in between the ln's

OpenStudy (anonymous):

Mhmmmm

OpenStudy (anonymous):

I think the ans is -1/169x^2-49 -1/(13x-7) 14(13x-7)

OpenStudy (anonymous):

Sorry delete the 14

OpenStudy (anonymous):

how did you get that??

OpenStudy (anonymous):

Let me double check

OpenStudy (anonymous):

You want to find out integral right?

OpenStudy (anonymous):

∫ 52/(x^2*sqrt(169x^2-49)) dx forget the 52 for now; we'll multiply the answer by 52 at the end ∫ 1/(x^2*sqrt(169x^2-49)) dx Let x= (7/13) sec t dx = (7/13) sec(t) tan(t) dt 169x^2-49 = 49 sec^2 t - 49 = 49(sec^2(t)-1) = 49 tan^2 t sqrt(169x^2-49) = 7 tan t ∫ 1/(x^2*sqrt(169x^2-49)) dx = ∫ (7/13) sec(t) tan(t) dt / (169/49) sec^2(t) (7 tan(t)) = (13/49) ∫ cos(t) dt (13/49) sin (t) Multiply by 52 (676/49) sin(t) + C (676/49) sqrt(1-cos^2(t)) + C sec(t) = (13/7) x cos(t) = (7/13x) cos^2(t) = 49/169x^2 sqrt(1-cos^2(t)) = sqrt(1-49/169x^2) = sqrt(169x^2-49) /13x (676/49) sqrt(1-cos^2(t)) = ((676/49)sqrt(169x^2-49) /13x = 52/ [49x sqrt(169x^2-49)] + C

OpenStudy (anonymous):

I hate math!

OpenStudy (anonymous):

ya. i gots it now. it was the damn chain rule. im having to use partial fractions to solve the problem. i had gotten an answer that was wrong but not too far off from what the real answer is. NOW i see whats up....with the chain rule i see how its 1/182 and a minus in between the ln's. yes my friend i hate math too. but with my major, ill never be done with it or physics. blarg.

OpenStudy (anonymous):

Well at least I tried

OpenStudy (anonymous):

you did. actually your answer made me realize it. so you did help me!

OpenStudy (anonymous):

I am glad it helped you

OpenStudy (anonymous):

so back to your question. im still looking for something that would help you. just so im clear, youre wanting to know how to find the empirical probability of a second generation plant correct?

OpenStudy (anonymous):

That's right

OpenStudy (anonymous):

I have to find find P(tall plants) and P(short plants)

OpenStudy (anonymous):

okay. im still looking for ya. :)

OpenStudy (anonymous):

I never seen a person tried so hard lol

OpenStudy (anonymous):

well i feel like if you give good things, good things come in return.

OpenStudy (anonymous):

I didn't give you nothing good though

OpenStudy (anonymous):

Nothing yet lol

OpenStudy (anonymous):

ya but you helped me see what my issue was. and im trying to prevent you jumping out your window. haha. its just been a long time since i had biology or done any probability probs. does your teacher want it in fractional form?

OpenStudy (anonymous):

bahaha

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

This chat so slow

OpenStudy (anonymous):

it really is give me a few mins. i found some info on probability. :)

OpenStudy (anonymous):

Gr8 Job

OpenStudy (anonymous):

I already jumped out my window

OpenStudy (anonymous):

okay so ive found something that is the only actual thing on probability. i couldnt find anything else that would explain probability other than this. this might help...heres the link. http://www.math.uah.edu/stat/prob/Probability.html about 3/4 of the page down, it talks about genetics (which from what i read im assuming youre doing) it talks about a pea plant and its parents. maybe this will help you! i hope!

OpenStudy (anonymous):

if you scroll down, youll see a in boldface "Genetics". thats where im referring to.

OpenStudy (anonymous):

You from N.Y

OpenStudy (anonymous):

I see it lol

OpenStudy (anonymous):

haha no im in Georgia.

OpenStudy (anonymous):

OO

OpenStudy (anonymous):

quite a ways away from NY. lol

OpenStudy (anonymous):

Pretty much lol

OpenStudy (anonymous):

im assuming thats where you are from. haha

OpenStudy (anonymous):

you know out of the millions of pages i searched through that was the only thing i found that even explained what probability is and how you can find it. everything else was just problems or articles on gambling. lolol

OpenStudy (anonymous):

You assumed it right

OpenStudy (anonymous):

Gambling wtf lol

OpenStudy (anonymous):

haha i know! lol

OpenStudy (anonymous):

gambling, cards and dice. go figure. nope my friend i dont even know what that is. hahahaha

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I bbroke my bonesq jumping out the window lol

OpenStudy (anonymous):

I still couldn't figure out my answer.

OpenStudy (anonymous):

but you can search for me on facebook....www.facebook.com/ashthedoll im sorry friend i tried. i fail with probability. lol

OpenStudy (anonymous):

I dont have facebook